OpenStudy (abbyrose0105):

Simplify the given expression to radical form and justify each step by identifying the properties of rational exponents used. All work must be shown. x to the two–thirds power, over x to the four–ninths power

OpenStudy (neer2890):

\[\frac{ x ^{\frac{ 2 }{ 3 }} }{ x ^{\frac{ 4 }{ 9 }} }=x ^{\frac{ 2 }{ 3 }}*x ^{-\frac{ 4 }{ 9 }}=x ^{(\frac{ 2 }{ 3 }-\frac{ 4 }{ 9 })}=?\]